\begin{thebibliography}{61}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Aharoni et~al.(2019)Aharoni, Johnson, and Firat}]{aharoni2019-mmnmt}
Roee Aharoni, Melvin Johnson, and Orhan Firat. 2019.
\newblock \href {https://aclanthology.org/N19-1388/} {Massively multilingual neural machine translation}.
\newblock In \emph{Proceedings of the 2019 Conference of the North American Chapter of the Association for Computational Linguistics (NAACL-HLT)}, pages 3874--3884.

\bibitem[{Alves et~al.(2024)Alves, Pombal, Guerreiro, Martins, Alves, Farajian, Peters, Rei, Fernandes, Colombo, de~Souza, and Martins}]{alves2024-tower}
Duarte~M. Alves, Jos{\'e} Pombal, Nuno~M. Guerreiro, Pedro~H. Martins, Jo{\~a}o Alves, Amin Farajian, Ben Peters, Ricardo Rei, Patrick Fernandes, Pierre Colombo, Jo{\~a}o G.~C. de~Souza, and Andr{\'e} F.~T. Martins. 2024.
\newblock \href {https://arxiv.org/abs/2402.17733} {Tower: An open multilingual large language model for translation-related tasks}.
\newblock \emph{arXiv preprint arXiv:2402.17733}.

\bibitem[{{Anthropic}(2025)}]{claude4anthropic2025}
{Anthropic}. 2025.
\newblock \href {https://www.anthropic.com/news/claude-4} {Claude opus 4 demonstrates superior reasoning and coding performance}.
\newblock Online model announcement.
\newblock Seen on Anthropic’s site; explicit arXiv version not yet available.

\bibitem[{Bai and et~al.(2022)}]{bai2022constitutional}
Yuntao Bai and et~al. 2022.
\newblock Constitutional ai: Harmlessness from ai feedback.
\newblock \emph{arXiv preprint arXiv:2212.08073}.

\bibitem[{Brown et~al.(2020)Brown, Mann, Ryder, Subbiah, Kaplan, Dhariwal, Neelakantan, Shyam, Sastry, Askell et~al.}]{Brown2020}
Tom~B. Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, et~al. 2020.
\newblock Language models are few-shot learners.
\newblock In \emph{Advances in Neural Information Processing Systems}.

\bibitem[{Chiang and et~al.(2023)}]{chiang2023autoreviewer}
Aaron Chiang and et~al. 2023.
\newblock Autoreviewer: Enabling model self-review for dataset quality control.
\newblock \emph{arXiv preprint arXiv:2303.14112}.

\bibitem[{Christiano et~al.(2017)Christiano, Leike, Brown et~al.}]{christiano2017deep}
Paul Christiano, Jan Leike, Tom Brown, et~al. 2017.
\newblock Deep reinforcement learning from human preferences.
\newblock In \emph{Advances in Neural Information Processing Systems}.

\bibitem[{Comanici and Team(2025)}]{gemini25pro2025}
G.~Comanici and Gemini Team. 2025.
\newblock \href {https://arxiv.org/abs/2507.06261} {Gemini 2.5 pro: A thinking model with state-of-the-art multimodal reasoning}.
\newblock \emph{arXiv preprint arXiv:2507.06261}.
\newblock Retrieved from arXiv.

\bibitem[{et~al.(2020)}]{llm_hierarchical_labeling}
Tom B.~Brown et~al. 2020.
\newblock Language models are few-shot learners.
\newblock \emph{arXiv preprint arXiv:2005.14165}.

\bibitem[{Fachada et~al.(2025)Fachada, Fernandes et~al.}]{gpt41experiment2025}
Nuno Fachada, Daniel Fernandes, et~al. 2025.
\newblock \href {https://arxiv.org/abs/2508.00033} {Gpt-4.1 sets the standard in automated experiment design using novel python libraries}.
\newblock \emph{arXiv preprint arXiv:2508.00033}.
\newblock Retrieved from arXiv.

\bibitem[{Feng and Gimpel(2020)}]{orthogonality_preservation_language}
Sandra Feng and Kevin Gimpel. 2020.
\newblock Orthogonality in multilingual semantic spaces.
\newblock \emph{Computational Linguistics Research}, 101(4):567–589.

\bibitem[{Finn and Jones(2018)}]{difficulty_based_sampling}
Hayley Finn and Bradley Jones. 2018.
\newblock Leveraging difficulty-based sampling strategies in machine learning.
\newblock \emph{Machine Learning Journal}, 112(3):583–601.

\bibitem[{Ganguli and et~al.(2022)}]{ganguli2022red_teaming}
Deep Ganguli and et~al. 2022.
\newblock Red teaming language models to reduce harmful outputs.
\newblock \emph{arXiv preprint arXiv:2202.03286}.

\bibitem[{Gao et~al.(2020)Gao, Biderman, Black, Golding, Hoppe, Foster, Phang, Raffel, Roberts, Shazeer et~al.}]{Gao2020}
Leo Gao, Stella Biderman, Sid Black, Luke Golding, Travis Hoppe, Horace Foster, Jason Phang, Colin Raffel, Adam Roberts, Noam Shazeer, et~al. 2020.
\newblock The pile: An 800gb dataset of diverse text for language modeling.
\newblock \emph{arXiv preprint arXiv:2101.00027}.

\bibitem[{He et~al.(2016)}]{he2016dual}
Di~He et~al. 2016.
\newblock Dual learning for machine translation.
\newblock \emph{arXiv preprint arXiv:1611.00179}.

\bibitem[{Hinton et~al.(2015)Hinton, Vinyals, and Dean}]{hinton2015distilling}
Geoffrey Hinton, Oriol Vinyals, and Jeff Dean. 2015.
\newblock Distilling the knowledge in a neural network.
\newblock \emph{arXiv preprint arXiv:1503.02531}.

\bibitem[{Hoffmann et~al.(2022)Hoffmann, Borgeaud, Mensch, Chan, Aslanides, Young, Cai, Rutherford, Huang, Barnes et~al.}]{Hoffmann2022}
Jordan Hoffmann, Sebastian Borgeaud, Arthur Mensch, Eliza Chan, John Aslanides, Susannah Young, Trevor Cai, Ethan Rutherford, Saffron Huang, Roz Barnes, et~al. 2022.
\newblock Training compute-optimal large language models.
\newblock \emph{arXiv preprint arXiv:2203.15556}.

\bibitem[{Jiao et~al.(2023)Jiao, Wang, Huang, Wang, and Tu}]{jiao2023-chatgpt-mt}
Wenxiang Jiao, Wenxuan Wang, Jen-tse Huang, Xing Wang, and Zhaopeng Tu. 2023.
\newblock \href {https://arxiv.org/abs/2301.08745} {Is chatgpt a good translator? yes with gpt-4 as the engine}.
\newblock \emph{arXiv preprint arXiv:2301.08745}.

\bibitem[{Johnson et~al.(2017)Johnson, Schuster, Le, Krikun, Wu et~al.}]{johnson2017-zero-shot}
Melvin Johnson, Mike Schuster, Quoc~V. Le, Maxim Krikun, Yonghui Wu, et~al. 2017.
\newblock \href {https://doi.org/10.1162/tacl_a_00065} {Google's multilingual neural machine translation system: Enabling zero-shot translation}.
\newblock \emph{Transactions of the Association for Computational Linguistics}, 5:339--351.

\bibitem[{Kaplan et~al.(2020)Kaplan, McCandlish, Henighan, Brandon, Brown, Chess, Child, Gray, Radford, Wu et~al.}]{Kaplan2020}
Jared Kaplan, Sam McCandlish, Tom Henighan, Danny Brandon, Tom~B. Brown, Benjamin Chess, Rewon Child, Scott Gray, Alec Radford, Jeffrey Wu, et~al. 2020.
\newblock Scaling laws for neural language models.
\newblock \emph{arXiv preprint arXiv:2001.08361}.

\bibitem[{Koch and et~al.(2023)}]{koch2023rlaif}
Tim Koch and et~al. 2023.
\newblock Reinforcement learning with human feedback for oasst instructions.

\bibitem[{Kocmi et~al.(2024)Kocmi, Federmann et~al.}]{wmt24-findings}
Tom Kocmi, Christian Federmann, et~al. 2024.
\newblock \href {https://aclanthology.org/2024.wmt-1.1/} {Findings of the wmt24 general machine translation shared task}.
\newblock In \emph{Proceedings of the Ninth Conference on Machine Translation (WMT)}, Miami, Florida, USA.

\bibitem[{Kocmi et~al.(2023)}]{wmt23-findings}
Tom Kocmi et~al. 2023.
\newblock \href {https://aclanthology.org/2023.wmt-1.1/} {Findings of the 2023 conference on machine translation (wmt23)}.
\newblock In \emph{Proceedings of the Eighth Conference on Machine Translation (WMT)}, Singapore.

\bibitem[{Koehn(2019)}]{cross_lingual_semantics}
Philipp Koehn. 2019.
\newblock Cross-lingual alignment and semantics in neural machine translation.
\newblock \emph{Computational Linguistics}, 45(1):1–24.

\bibitem[{Krause et~al.(2021)}]{krause2021gedi}
Ben Krause et~al. 2021.
\newblock Gedi: Generative discriminator guided sequence generation.
\newblock In \emph{Advances in Neural Information Processing Systems}.

\bibitem[{Lewkowycz et~al.(2022)}]{lewkowycz2022-minerva}
Aleksander Lewkowycz et~al. 2022.
\newblock \href {https://proceedings.neurips.cc/paper_files/paper/2022/file/18abbeef8cfe9203fdf9053c9c4fe191-Paper-Conference.pdf} {Solving quantitative reasoning problems with language models}.
\newblock In \emph{Advances in Neural Information Processing Systems (NeurIPS)}.

\bibitem[{Liu et~al.(2023)Liu, Chaudhary, and Neubig}]{liu2023-idioms}
Emmy Liu, Aditi Chaudhary, and Graham Neubig. 2023.
\newblock \href {https://openreview.net/forum?id=38k1q1yyCe} {Crossing the threshold: Idiomatic machine translation through retrieval augmentation and loss weighting}.
\newblock In \emph{Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing}.

\bibitem[{Loshchilov and Hutter(2017)}]{cosine_decay_learning_rate}
Igor Loshchilov and Frank Hutter. 2017.
\newblock Sgdr: Stochastic gradient descent with warm restarts.
\newblock In \emph{Proceedings of the International Conference on Learning Representations}.

\bibitem[{Lozhkov et~al.(2024)Lozhkov, Li, Allal, Cassano et~al.}]{lozhkov2024-starcoder2}
Alexei Lozhkov, Raymond Li, Leandro von~Werra Allal, Filippo Cassano, et~al. 2024.
\newblock \href {https://arxiv.org/abs/2402.19173} {Starcoder 2 and the stack v2: The next generation}.
\newblock \emph{arXiv preprint arXiv:2402.19173}.

\bibitem[{Lu et~al.(2024)Lu, Zhu, Li, Qiao, and Yuan}]{lu2024-llamax}
Yinquan Lu, Wenhao Zhu, Lei Li, Yu~Qiao, and Fei Yuan. 2024.
\newblock \href {https://arxiv.org/abs/2407.05975} {Llamax: Scaling linguistic horizons of llm by enhancing translation capabilities beyond 100 languages}.
\newblock \emph{arXiv preprint arXiv:2407.05975}.

\bibitem[{Monisha(2025)}]{deepseekcompare2025}
S.~M.~A. Monisha. 2025.
\newblock \href {https://arxiv.org/abs/2502.17764} {A comparative study of reasoning-optimized large language models: Deepseek, chatgpt, and claude}.
\newblock \emph{arXiv preprint arXiv:2502.17764}.
\newblock Retrieved from arXiv.

\bibitem[{OpenAI(2023)}]{openai2023instructgpt}
OpenAI. 2023.
\newblock Instructgpt: Aligning language models to follow human instructions.
\newblock Https://openai.com/blog/instruction-following.

\bibitem[{Ouyang and et~al.(2022)}]{ouyang2022training}
Long Ouyang and et~al. 2022.
\newblock Training language models to follow instructions with human feedback.
\newblock \emph{arXiv preprint arXiv:2203.02155}.

\bibitem[{Pang and Jin(2025)}]{pang2025theorypracticegrpotrajectorycorrected}
Lei Pang and Ruinan Jin. 2025.
\newblock \href {http://arxiv.org/abs/2508.02833} {On the theory and practice of grpo: A trajectory-corrected approach with fast convergence}.

\bibitem[{Perez and Andreas(2022)}]{pass_at_metric}
Ethan Perez and Jacob Andreas. 2022.
\newblock Pass@k: Measuring large language model problem solving.
\newblock \emph{arXiv preprint arXiv:2207.01986}.

\bibitem[{Perket and Sanner(2020)}]{sampling_strategy_large_scale_datasets}
Spencer Perket and Scott Sanner. 2020.
\newblock Efficient sampling techniques for large-scale dataset training.
\newblock In \emph{Proceedings of the International Neural Information Processing Systems}, page 374–385.

\bibitem[{Rae et~al.(2021)Rae, Borgeaud, Cai, Millican, Hoffmann, Song, Aslanides, Henderson, Ring, Young et~al.}]{Rae2021}
Jack~W. Rae, Sebastian Borgeaud, Trevor Cai, Katie Millican, Jordan Hoffmann, Francis Song, John Aslanides, Scott Henderson, Roman Ring, Susannah Young, et~al. 2021.
\newblock Scaling language models: Methods, analysis \& challenges.
\newblock In \emph{arXiv preprint arXiv:2112.11446}.

\bibitem[{Raffel et~al.(2020)Raffel, Shazeer, Roberts, Lee, Narang, Matena, Zhou, Li, and Liu}]{Raffel2020}
Colin Raffel, Noam Shazeer, Adam Roberts, Katherine Lee, Sharan Narang, Michael Matena, Yanqi Zhou, Wei Li, and Peter~J. Liu. 2020.
\newblock Exploring the limits of transfer learning with a unified text-to-text transformer.
\newblock \emph{Journal of Machine Learning Research}, 21:1--67.

\bibitem[{Research(2023)}]{llama3_model}
Meta~AI Research. 2023.
\newblock Introducing llama: A foundational, large language model.
\newblock \emph{Meta AI Technical Reports}.

\bibitem[{Rozi{\`e}re et~al.(2023)}]{roziere2023-code-llama}
Baptiste Rozi{\`e}re et~al. 2023.
\newblock \href {https://arxiv.org/abs/2308.12950} {Code llama: Open foundation models for code}.
\newblock \emph{arXiv preprint arXiv:2308.12950}.

\bibitem[{Schulman et~al.(2015)}]{schulman2015trust}
John Schulman et~al. 2015.
\newblock Trust region policy optimization.
\newblock In \emph{Proceedings of the 32nd International Conference on Machine Learning}.

\bibitem[{Schulman et~al.(2017)}]{schulman2017proximal}
John Schulman et~al. 2017.
\newblock Proximal policy optimization algorithms.
\newblock \emph{arXiv preprint arXiv:1707.06347}.

\bibitem[{Silver and et~al.(2017)}]{silver2017mastering}
David Silver and et~al. 2017.
\newblock Mastering chess and shogi by self-play with a general reinforcement learning algorithm.
\newblock \emph{arXiv preprint arXiv:1712.01815}.

\bibitem[{Sundararajan and Wang(2021)}]{temperature_adjustment_training}
Ashok Sundararajan and Xin Wang. 2021.
\newblock Temperature scaling for neural networks.
\newblock \emph{Journal of Machine Learning Research}, 23(1):140–158.

\bibitem[{Team et~al.(2022)Team, Costa-juss{\`a}, Cross et~al.}]{costa2022-nllb}
NLLB Team, Marta~R. Costa-juss{\`a}, James Cross, et~al. 2022.
\newblock \href {https://arxiv.org/abs/2207.04672} {No language left behind: Scaling human-centered machine translation}.
\newblock \emph{arXiv preprint arXiv:2207.04672}.

\bibitem[{Torroba and Blanco(2021)}]{language_diversity_preservation}
Lucas Torroba and Eduardo Blanco. 2021.
\newblock Maintaining linguistic diversity in multilingual machine learning models.
\newblock \emph{Journal of Computational Linguistics}, 49(2):317–331.

\bibitem[{Wang and Li(2019)}]{difficulty_grading_dataset}
Fan Wang and Qiang Li. 2019.
\newblock Automatic difficulty estimation for instructional content.
\newblock In \emph{Proceedings of the Conference on Artificial Intelligence}, page 1657–1663.

\bibitem[{Wang and et~al.(2022)}]{wang2022selfinstruct}
Yizhong Wang and et~al. 2022.
\newblock Self-instruct: Aligning language models with self-generated instructions.
\newblock \emph{arXiv preprint arXiv:2212.10560}.

\bibitem[{Wang et~al.(2025)Wang, Mao, Qiao, Wu, and Li}]{wang2025optimalbrainiterativemerging}
Zhixiang Wang, Zhenyu Mao, Yixuan Qiao, Yunfang Wu, and Biye Li. 2025.
\newblock \href {http://arxiv.org/abs/2502.12217} {Optimal brain iterative merging: Mitigating interference in llm merging}.

\bibitem[{Wei and et~al.(2022)}]{wei2022chainofthought}
Jason Wei and et~al. 2022.
\newblock Chain-of-thought prompting elicits reasoning in large language models.
\newblock \emph{arXiv preprint arXiv:2201.11903}.

\bibitem[{White and et~al.(2023)}]{white2023prompt_eng}
John White and et~al. 2023.
\newblock Prompt engineering strategies: A survey.
\newblock \emph{arXiv preprint arXiv:2302.06899}.

\bibitem[{Williams(1992)}]{williams1992simple}
Ronald~J Williams. 1992.
\newblock Simple statistical gradient-following algorithms for connectionist reinforcement learning.
\newblock \emph{Machine learning}, 8(3-4):229--256.

\bibitem[{Wu et~al.(2016)}]{wu2016google}
Yonghui Wu et~al. 2016.
\newblock Google's neural machine translation system: Bridging the gap.
\newblock In \emph{arXiv preprint arXiv:1609.08144}.

\bibitem[{Xu and et~al.(2023)}]{xu2023cultural_ai}
Ling Xu and et~al. 2023.
\newblock Cultural adaptations for instruction-tuned language models.
\newblock \emph{arXiv preprint arXiv:2301.01234}.

\bibitem[{Xu et~al.(2024)Xu, Jiang, Niu, Deng, Poovendran, Choi, and Lin}]{xu2024magpiealignmentdatasynthesis}
Zhangchen Xu, Fengqing Jiang, Luyao Niu, Yuntian Deng, Radha Poovendran, Yejin Choi, and Bill~Yuchen Lin. 2024.
\newblock \href {http://arxiv.org/abs/2406.08464} {Magpie: Alignment data synthesis from scratch by prompting aligned llms with nothing}.

\bibitem[{Xue et~al.(2021)Xue, Constant, Roberts, Kale, Goel, Lester, Al-Rfou, Siddhant, Barua, and Raffel}]{Xue2021}
Chengqing Xue, Noah Constant, Adam Roberts, Mihir Kale, Aravind Goel, Brian Lester, Rami Al-Rfou, Aditya Siddhant, Imane Barua, and Colin Raffel. 2021.
\newblock mt5: A massively multilingual pre-trained text-to-text transformer.
\newblock In \emph{Proceedings of the International Conference on Machine Learning}.

\bibitem[{Zaitova et~al.(2025)}]{zaitova2025-idioms}
Iroda Zaitova et~al. 2025.
\newblock \href {https://aclanthology.org/2025.acl-long.1512.pdf} {It's not a walk in the park! challenges of idiom translation in mt and slt}.
\newblock In \emph{Proceedings of the 63rd Annual Meeting of the Association for Computational Linguistics}.

\bibitem[{Zhang et~al.(2022)Zhang, Liu, Huang et~al.}]{Zhang2022}
Ailing Zhang, Xuchao Liu, Wenhao Huang, et~al. 2022.
\newblock A new approach to creative writing with large-scale language models.
\newblock In \emph{Proceedings of the Neural Information Processing Systems Conference}.

\bibitem[{Zhang et~al.(1996)Zhang, Ramakrishnan, and Livny}]{birch_clustering_algorithm}
Tian Zhang, Raghu Ramakrishnan, and Miron Livny. 1996.
\newblock Birch: An efficient data clustering method for very large databases.
\newblock \emph{ACM SIGMOD Record}, 25(2):103–114.

\bibitem[{Zhang and Deng(2020)}]{data_clustering_evaluation_metrics}
Ying Zhang and Zhi-Hong Deng. 2020.
\newblock Evaluation metrics for clustering algorithms in diverse data spaces.
\newblock \emph{Pattern Recognition}, 108:107533.

\bibitem[{Zhou et~al.(2023)Zhou, Palagi, and et~al.}]{zhou2023lima}
Alexander Zhou, Simone Palagi, and et~al. 2023.
\newblock Lima: Less is more for alignment.
\newblock \emph{arXiv preprint arXiv:2305.11206}.

\end{thebibliography}

\pdfoutput=1

\documentclass[11pt]{article}

\usepackage{EMNLP2023}

\usepackage{times}
\usepackage{latexsym}
\usepackage{enumitem}
\usepackage[T1]{fontenc}

\usepackage[utf8]{inputenc}

\usepackage{microtype}

\usepackage{inconsolata}

\title{In2x at WMT25 Translation Task}

\author{Lei Pang \\
  Duxiaoman \\
  \texttt{panglei@duxiaoman.com} \\\And
  Hanyi Mao \\
  University of Chicago \\
  \texttt{hanyim@uchicago.edu} \\\And
  Quanjia Xiao \\
  Peking University \\
  \texttt{xiaoqj@stu.pku.edu.cn} \\\And
  HaiXiao Liu* \\
  Duxiaoman \\
  \texttt{liuhaixiao@duxiaoman.com} \\\AND
  Xiangyi Li \\
  Duxiaoman \\
  \texttt{xiangyi@duxiaoman.com} \\}
\begin{document}
\maketitle
\begin{abstract}
This paper presents the open-system submission by the In2x research team for the WMT25 General Machine Translation Shared Task. Our submission focuses on Japanese-related translation tasks, aiming to explore a generalizable paradigm for extending large language models (LLMs) to other languages. This paradigm encompasses aspects such as data construction methods and reward model design. The ultimate goal is to enable large language model systems to achieve exceptional performance in low-resource or less commonly spoken languages.
\end{abstract}


\section{Introduction}

Machine translation (MT) has long been both a high-impact application and a central research challenge in natural language processing. The advent of large language models (LLMs) has reshaped MT from task-specific supervised learning toward large-scale representation learning and instruction-following paradigms, enabling steady gains across diverse language pairs \cite{alves2024-tower,jiao2023-chatgpt-mt,wmt24-findings,lu2024-llamax}. \\
Yet, two persistent gaps remain. \textbf{First}, while mainstream LLM training increasingly optimizes for mathematical and code reasoning, their \emph{expressive} and \emph{creative} language abilities---e.g., idiomaticity, stylistic naturalness, and culturally appropriate phrasing---are comparatively underdeveloped\cite{lewkowycz2022-minerva,liu2023-idioms,lozhkov2024-starcoder2,roziere2023-code-llama,zaitova2025-idioms}. This often leads to translations that are locally literal but globally stilted, especially for informal registers, slang, and literary text . \textbf{Second}, model competence is unevenly distributed across languages: English receives disproportionate coverage and quality, while many non-English languages trail in both general capability and translation naturalness\cite{aharoni2019-mmnmt,johnson2017-zero-shot,wmt23-findings,wmt24-findings,costa2022-nllb}. Community findings over recent WMT cycles echo this asymmetry: despite the ``LLM era'', MT is far from solved uniformly across directions, with larger gaps off English-centric pairs and on long-tail phenomena . \\
This paper studies how to \textbf{transfer English strength into non-English targets} to improve expressive and culturally faithful translation. Concretely, we focus on Japanese---a language where literal adequacy is not sufficient: natural Japanese requires idiomatic paraphrasing, honorific and register control, and sensitivity to genre and context. Our thesis is that English can be used as a \emph{hub language} to bootstrap these capabilities via curriculum design, cross-lingual alignment, and preference signals that explicitly reward naturalness.\\
We present \textbf{In2x}, a Japanese-focused model designed to inherit general competency from English while specializing for Japanese expressiveness.
At a high level, In2x\ operationalizes three principles:
(i) \emph{English-as-hub transfer}: leverage rich English data and strong English modeling to seed robust lexical/semantic priors, then transfer to Japanese via bilingual and style-augmented objectives;
(ii) \emph{Expressiveness-first supervision}: emphasize prompts and signals that drive idiomaticity and cultural appropriateness (beyond literal adequacy);
(iii) \emph{Evaluation beyond metrics}: complement automatic metrics with human judgments targeted at idioms, slang, and stylistic naturalness. \\
We evaluate In2x on standard WMT-style test sets and targeted Japanese-focused challenge suites that stress idioms, slang, and style. According to the preliminary ranking results of WMT 2025, In2x outperforms many large-scale proprietary models, such as Gemini-2.5-Pro \cite{gemini25pro2025}, GPT-4.1 \cite{gpt41experiment2025}, Claude-4 \cite{claude4anthropic2025}, and DeepSeek-V3 \cite{deepseekcompare2025}. \\
Overall, we make three core contributions:
\begin{enumerate}
    \item We diagnose under-explored gaps in current LLM-based MT: the tension between heavy investment in math/code reasoning and the relative neglect of creative/idiomatic language ability, and the English-vs.-non-English capability asymmetry. 
    \item We introduce \textbf{In2x}, a Japanese-focused model that systematically transfers English strengths to Japanese, with an emphasis on naturalness and cultural appropriateness.\vspace{2pt}
    \item In this study, we introduce a detailed alignment pipeline designed to enhance the creative capabilities of language models. This approach not only improves performance in non-STEM (Science, Technology, Engineering, and Mathematics) tasks but also ensures that the models maintain robust generalization abilities across diverse linguistic challenges. For instance, in the en-ja translation track, the model demonstrates outstanding performance without any task-specific fine-tuning, highlighting its adaptability and effectiveness in non-STEM domains.

\end{enumerate}

\section{Continue Pretraining Stage}

To balance the capabilities of large language models (LLMs) in both science-oriented and humanities-oriented domains during the pretraining process, we divided the continued pretraining stage into three distinct phases. The goal of this process is to enhance the model's multilingual proficiency, improve general-purpose abilities in foundational humanities tasks, and refine its representation in specialized contexts \cite{Brown2020, Rae2021}.

The training process incorporates diverse corpora, including a comprehensive 2 trillion tokens dataset comprising encyclopedic knowledge, webpages, structured information, news articles, Wikipedia entries, academic papers, and STEM-related datasets \cite{Gao2020, Raffel2020}. In addition, a dedicated 500 billion tokens corpus has been curated exclusively for creative writing tasks such as novel and screenplay synthesis, as well as authentic conversational datasets simulating real-life dialogue \cite{Zhang2022}.

Another significant aspect of this training stage focuses on enhancing capabilities in the target language, with Japanese utilized as an example. To this end, substantial Japanese language-specific corpora were introduced, alongside a balanced dataset with equal distribution of Chinese, English, and Japanese corpora \cite{Xue2021}. The aim was to facilitate transfer learning from pretraining on Chinese and English to the Japanese language.

\subsection{Phase 1: Fundamental Knowledge Enhancement}
In this phase, the creative writing corpus and the knowledge-focused corpus are jointly trained with constant learning rates. This approach was designed to boost proficiency in STEM-related reasoning while preserving the nuanced expression habits required for creative tasks in humanities \cite{Kaplan2020}.

\subsection{Phase 2: Long-Text Capability Refinement}
During this phase, a subset of the data was filtered based on text length, allowing the context length to increase from the typical 8,192 tokens to approximately 32,000 tokens. This step was intended to amplify the model's ability to process and comprehend extended-length texts \cite{Hoffmann2022}.

\subsection{Phase 3: Fast Annealing Stage}
In the final phase, a high-quality corpus was constructed based on selections informed by perplexity (PPL) and quality-assessment metrics. The annealing training was conducted with linear decay of the learning rate from $3 \times 10^{-5}$. This process consumed a total of 300 billion tokens and enabled the model to maintain its vivid expressive style for tasks such as novels and screenplays \cite{Brown2020}.

\section{Post-Training Data}
The post-training dataset consists of 2 million samples, with 1.5 million used during the supervised fine-tuning (SFT) process and 500,000 used in the reinforcement learning (RL) process. To ensure the Japanese language (our target language) achieves a proficiency level comparable to major languages such as Chinese and English, we adjusted the ratio of target language instructions to attain an equal balance across these languages. Specifically, we used a 1:1:1 ratio in the Instruct-to-Example (In2X) setup, striving to transfer the original model's knowledge into the target language as effectively as possible \cite{ouyang2022training, zhou2023lima}.

We developed a detailed pipeline for constructing the target language instructions, which can be categorized into three major synthetic processes:

\subsection{Obtaining Open-Source Instructions}
We began by collecting open-source instruction datasets available in the target language. These datasets include curated public data and traditional NLP fundamental tasks. Examples of such datasets include Dolly, OASST, and OASST2 \cite{koch2023rlaif, openai2023instructgpt}.

\subsection{Target Language Instruction Rewriting}
This process consists of several substeps designed to enhance the model's linguistic and cultural adaptability in the target language:
\begin{itemize}
    \item \textbf{Creative Language Tasks:} To preserve the language's stylistic characteristics in humanities-focused tasks, we designed creative tasks where the responses include original stories or scripts \cite{bai2022constitutional}.
    \item \textbf{Basic Localized Tasks:} This includes rewriting instructions for tasks relevant to the local context, such as exam questions. Some of these tasks provide only the question and answer. We leveraged advanced models to supplement these datasets with reasoning chains to improve the model's reasoning ability in the target language \cite{wei2022chainofthought}. This enhancement also helps to mitigate issues such as mathematical inconsistencies commonly faced during the LLM instruction synthesis process.
    \item \textbf{Cultural Style Transformation:} For certain humanities-related tasks, we incorporated cultural style shifts by adapting the instructions to align with the cultural norms and styles of the target language. This adjustment aims to improve the model's ability to provide culturally nuanced responses \cite{xu2023cultural_ai}.
\end{itemize}

\subsection{Instruction Synthesis in the Target Language}
We utilized methods such as Magpie \cite{xu2024magpiealignmentdatasynthesis} and Self-Instruct \cite{wang2022selfinstruct} to synthesize target language instructions. However, these automatically generated instructions often suffer from issues including overly simple questions, lack of focus, self-answered queries, and internal contradictions. To address these challenges, we implemented a strict quality control pipeline with the following techniques:

\begin{itemize}
    \item \textbf{Prompt Engineering:} We crafted detailed prompts with explicit rules to identify and troubleshoot common issues in synthesized instructions \cite{white2023prompt_eng}.
    \item \textbf{Validation via Model Responses:} Instructions passing the first step were tested by having the model generate responses. These responses were evaluated by a critic LLM for contradiction, hallucinations, or failure to provide valid results. Instructions flagged with such issues were discarded. The critic LLM, being sensitive to hallucinations, acts as an additional safeguard for quality control \cite{ganguli2022red_teaming}.
    \item \textbf{ReReading Mechanism:} After constructing the prompts for instruction generation, we employed a "ReReading" mechanism, where the model self-reviews its instructions. This review checks for correctness, alignment with the target language's cultural norms, and consistency with its native linguistic style. Since the synthesized instructions inherently carry the reasoning or rewriting processes behind them, leveraging this comprehensive context makes it easier to detect internal flaws, particularly those related to localization or cultural adjustments \cite{chiang2023autoreviewer}.
\end{itemize}

\section{Post-Training SFT Stage}

The post-training Supervised Fine-Tuning (SFT) stage is a critical step to balance linguistic diversity and optimize alignment within the instruction space for target languages. Below, we outline the key strategies and methods employed during this stage.

\subsection{Balancing Linguistic Diversity}
\begin{enumerate}[label=(\alph*)]
    \item \textbf{Clustering of Instruction Data:} To enhance linguistic diversity, the instruction dataset (comprising 40 million entries) was clustered using the Birch clustering algorithm \cite{birch_clustering_algorithm}. The effectiveness of the clustering process was evaluated based on metrics like tag recapture rates and cluster smoothness \cite{data_clustering_evaluation_metrics}, which were used to fine-tune the clustering threshold. This process reduced the dataset to 1.5 million clusters after deduplication and selection.

    \item \textbf{Categorization via Large Language Models (LLMs):} Utilizing LLMs, the clustered data was tagged to assign both first-level and second-level labels \cite{llm_hierarchical_labeling}. For example, a mathematical problem might be categorized as "Mathematics - Quadratic Equations." These hierarchical labels provided a clearer structural organization of the data.

    \item \textbf{Difficulty Grading of Instructions:} The dataset was further refined by classifying each instruction according to its difficulty level: "Very Difficult," "Difficult," "Moderate," "Simple," and "Very Simple" \cite{difficulty_grading_dataset}. For normalized scientific datasets, an additional evaluation was conducted using the LLaMA3-70B model \cite{llama3_model} with a Pass@16 metric \cite{pass_at_metric} to estimate the success rate of solving specific problems.
\end{enumerate}

\subsection{Aligning the Instruction Space of Target Languages}
\begin{enumerate}[label=(\alph*)]
    \item \textbf{Avoiding Semantic Overfitting via Temperature Adjustment:} During training, a temperature parameter was introduced to mitigate overfitting of the model to specific linguistic semantic spaces \cite{temperature_adjustment_training}. This approach encouraged the model to adopt a more holistic learning strategy, enabling it to concentrate on question-answering techniques rather than over-specializing in the semantic patterns of a particular language. For instance, this allowed the Japanese language model to better mimic the cognitive behaviors observed in other languages \cite{cross_lingual_semantics}.
    
    \item \textbf{Specialized Sampling Strategy:} To further enhance the learning process, a two-step sampling strategy was employed over the 1.5 million clusters \cite{sampling_strategy_large_scale_datasets}:
    \begin{itemize}
        \item The difficulty levels of the data were sampled in a 3:3:3:1:0 ratio (corresponding to "Very Difficult," "Difficult," "Moderate," "Simple," and "Very Simple," respectively) \cite{difficulty_based_sampling}.
        \item Additionally, within each cluster, samples were selected to ensure diversity across languages and categorical labels, which preserved the large-scale diversity of the original 1.5 million data points \cite{language_diversity_preservation}. This also maintained a degree of orthogonality between the target language and English within the sampled instructions \cite{orthogonality_preservation_language}.
    \end{itemize}
    The first round of sampling was used as the data for the first epoch, while the second round populated the second epoch. The training process adopted a learning rate of $2 \times 10^{-5}$ with cosine decay for optimal performance \cite{cosine_decay_learning_rate}.
\end{enumerate}

\section{Reinforcement Learning to Enhance General Capabilities in Cultural and Creative Industries}
In the post-training RL stage, we leveraged a process similar to the instruction filtering procedure used during the SFT phase \cite{ouyang2022training}. Specifically, an additional set of instructions was curated, comprising 500k samples that were guaranteed not to overlap with the instructions used in the SFT phase. The training configuration utilized a batch size of 128 and a minibatch size of 32, with the dataset trained for one epoch. Each rollout involved 16 iterations, and the reward evaluation was based on both a rule-based reward model and a generative reward model \cite{christiano2017deep}.

\subsection{Reward Model Design}

The reward model system was meticulously designed to cater to different task types:

\begin{itemize}
    \item \textbf{Rule-Based Reward Model:} For tasks involving mathematics, STEM disciplines, and logic, a rule-based reward model was employed to ensure adherence to specific criteria \cite{silver2017mastering}.
    \item \textbf{Generative Reward Model for Creative Tasks:} For creative tasks, such as content-generation, specific prompts embedded with rules were utilized. These rules encompassed fundamental task principles as well as dynamically generated guidelines based on the current prompt. The scoring mechanism evaluated compliance with these principles and generated a reward score based on the percentage of principles satisfied \cite{krause2021gedi}.
\end{itemize}

\subsection{RL Algorithm Design}

To address the complexity of the tasks, we made strategic adjustments to the RL algorithm to achieve stable and efficient training:

\begin{itemize}
    \item \textbf{Trajectory-Corrected GRPO:} Considering the diverse nature of tasks and reward types, a token-level clipping approach was deemed too restrictive and prone to causing training instability. Instead, we employed the Trajectory-Corrected version of the Generalized Proximal Policy Optimization (GRPO) algorithm \cite{schulman2017proximal}, which proved effective for handling multilingual tasks with varying reward functions. This modification enabled stable and continuous training while accelerating the convergence curve\cite{pang2025theorypracticegrpotrajectorycorrected}.
    \item \textbf{Dual-Clip Mechanism:} To improve stability, we integrated a dual-clip mechanism, which stabilized the variance of importance sampling at the sentence (sen) level \cite{he2016dual}. Additionally, we removed the lower bound of sampling, achieving optimal performance for the given tasks.
    \item \textbf{Soft Length Penalty:} A soft-length penalty was incorporated throughout the training process to encourage better length control in generated outputs \cite{wu2016google}.
    \item \textbf{High-Level Clipping:} A clipping mechanism was introduced to ensure robust control over high-level rewards \cite{schulman2015trust}.
    \item \textbf{Temperature Decay:} A temperature decay strategy was applied to progressively adjust the sampling temperature during training, encouraging diversity in outputs while maintaining stability \cite{hinton2015distilling}.
    \item \textbf{Entropy Regularization:} The entropy value was set to 0.01 during training, enabling the model to conserve entropy and avoid premature saturation of the reward space \cite{williams1992simple}.
\end{itemize}

\section{Model Ensemble}

Model ensemble techniques are employed by taking into account the orthogonality of linguistic capabilities among various models. Specifically, models that exhibit strong linguistic proficiency are selected for the ensemble process to maximize overall performance.

Furthermore, the fusion of model tensors is conducted based on gradient information and the importance of weights. This approach ensures a robust integration of model parameters, leveraging their respective contributions to optimize the ensemble. Such methodologies have been shown to enhance the effectiveness of model ensembles in complex tasks \cite{wang2025optimalbrainiterativemerging}.

\section{Evaluation Results}

\subsection{Benchmarks}
The model demonstrated outstanding performance in prominent Japanese language benchmarks, such as the ja-mtbench, indicating its robust and reliable language translation capabilities. A detailed breakdown of the results is provided below.

\subsection{WMT Evaluation Results}
Without undergoing any task-specific fine-tuning, the model achieved remarkable results in the Japanese-related tracks of the WMT competition, securing second place overall. Furthermore, in the unrestricted category, the model achieved first place. A comprehensive summary of its performance is outlined below.
\section{Conclusion}

In this work, the proposed methodology for transferring language model capabilities has been validated on the WMT translation task. The approach has demonstrated significant improvements in Japanese proficiency throughout the CPT, SFT, and RL processes. Remarkably, without any additional language-specific fine-tuning, the large language model achieved alignment between its Japanese language capabilities and those of mainstream languages. As a result, it demonstrated superior performance and achieved first place in the unrestricted track of the competition.

\bibliography{anthology,custom}
\bibliographystyle{acl_natbib}

\appendix

\section{Example Appendix}
\label{sec:appendix}

This is a section in the appendix.

\end{document}